OpenStudy (anonymous):

When taking a derivative of (1/2)^x is it x(1/2)^(x-1) or ln(1/2)u'

OpenStudy (anonymous):

Are the two the same?

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

the "power rule" is used if the bases is the variable, and the exponent is a number

OpenStudy (anonymous):

\[\frac{d}{dx}b^x=b^x\ln(b)\]